OpenStudy (abdullah1995):

A snowball is melting in the sun. It is noted that its surface decreases at a rate of 3cm2/s at the moment when its diameter is 8/picm. The goal here is to determine the rate at which the diameter varies at that same moment. To solve this problem, let x be the diameter of the snowball in cm, A its area in cm2, and t the time in seconds (s). (a) Express A as a function of x. A= __________ cm2 (b) What is the value of dA over dx when x=8/pi? Give the exact value. dA over dx= __________ cm (c) Using our previous results, give the (exact) value of dx over dt when x=8/pi cm.

OpenStudy (amistre64):

seems like we want to use the formula for the surface area of a sphere.

OpenStudy (abdullah1995):

well i need help ASAP pls

OpenStudy (amistre64):

that is the help .... what is the formula for the surface area of a sphere?

OpenStudy (abdullah1995):

i can find part A but i am stuck on part B

OpenStudy (amistre64):

we take an implicit derivative of the Area formula

OpenStudy (amistre64):

what did you get for part a?

OpenStudy (abdullah1995):

part b is dA/dx

OpenStudy (abdullah1995):

4pi(x/2)^2 for part a

OpenStudy (amistre64):

4 pi r^2 but r = d/2 4 pi (d/2)^2 simplify 4 pi d^2/4 = pi d^2 but they want x instead of d ... like a name matters sooo A = pi x^2

OpenStudy (abdullah1995):

is that the answer for A ^ ?

OpenStudy (amistre64):

i hope so, otherwise im just doing my taxes :)

OpenStudy (abdullah1995):

ok so whats the part B then ? C:

OpenStudy (amistre64):

part B is to find the derivative wrt.x of course ... dA/dx = ?

OpenStudy (abdullah1995):

pls give me the answer i am confused T_T

OpenStudy (amistre64):

idont just give out answers, that is NOT what this site is used for.

OpenStudy (abdullah1995):

ok if i post my work here can you help me identify the error ?

OpenStudy (amistre64):

ye si can

OpenStudy (abdullah1995):

ok i will post my solution 1 min pls

OpenStudy (abdullah1995):

(b) is asking for dA/dx so dA/dx = dA/dt * dt/dx that's all i know. am i on the right path ?

OpenStudy (amistre64):

dA/dx is just a 'normal' derivative. in this case its a simple power rule. given A = pi x^2, just read this as if its was something like: y = 3x^2

OpenStudy (abdullah1995):

what do you mean by normal derivative ? am i just suppose to plug 8/pi in A=pi x^2 to find the answer for B ?

OpenStudy (amistre64):

What is the value of dA/dx when x=8π? well, what is the derivative of A = pi x^2? derive, and then plug in x=8pi

OpenStudy (amistre64):

we have a constant and x to a power ... use the power rule

OpenStudy (abdullah1995):

so the derivative is pi*2x

OpenStudy (amistre64):

refrech with f5 to clear up the odd marks

OpenStudy (amistre64):

yes, the derivative of A = pi x^2 is A' = 2pi x

OpenStudy (amistre64):

your post says x=8pi (not x=8/pi) ... so if that is an error, then use what your problem wants you to use.

OpenStudy (abdullah1995):

yes i just copied the question so it is not showing the divide sign

OpenStudy (amistre64):

but yes, IF x=8/pi the dA/dx = 16

OpenStudy (amistre64):

and part C you alluded to earlier, the chain rule :\[\frac{dA}{dt}=\frac{dA}{dx}\frac{dx}{dt}\]

OpenStudy (abdullah1995):

ok i just fixed my question up i hope that clarifies things up

OpenStudy (amistre64):

oh it does :)

OpenStudy (abdullah1995):

so everything we were doing so far is correct ?

OpenStudy (amistre64):

yes, everything so far has been fine

OpenStudy (amistre64):

and since we know dA/dt from the infomation, and dA/dx from part B, the rest is just simple division

OpenStudy (abdullah1995):

ok so i understand part a and b now but i am still confused on part C

OpenStudy (amistre64):

\[\frac{dA}{dt}=\frac{dA}{dx}\frac{dx}{dt}\] we know everything but dx/dt ...

OpenStudy (abdullah1995):

ohhhhh ok so it's -3 / 16 ?

OpenStudy (amistre64):

yes :)

OpenStudy (abdullah1995):

alright thanks for the help. i will try to enter these answers and if they are wrong i will be back again for more help c: